Math teachers frequently claim that doing hw helps students get better grades in their math classes. To test this claim, a survey of high school math students was conducted from which the following results were obtained:

- 48% completed math hw regularly

- 55% have a B average or better in math class

- 40% do not complete math hw regularly AND have less than a B average in math class

Do these data support the claim that students who complete math hw regularly are more likely to get a B or better in math class? Justify your answer.
Answer #1

completed math hw regularly = 48%

have a B average or better in math class = 55%

do not complete math hw regularly AND have less than a B average in math class = 40%

complete math hw regularly OR have more than a B average in math class = 100 - do not complete math hw regularly AND have less than a B average in math class

complete math hw regularly OR have more than a B average in math class = 100 - 40

complete math hw regularly OR have more than a B average in math class = 60%

Now,

complete math hw regularly OR have more than a B average in math class = completed math hw regularly +  have a B average or better in math class - complete math hw regularly AND have more than a B average in math class

60 = 48 + 55 -  complete math hw regularly AND have more than a B average in math class

complete math hw regularly AND have more than a B average in math class = 48 + 55 - 60

complete math hw regularly AND have more than a B average in math class = 43%

Only 43% of the students who has done their hw regularly has scored more than a B average in maths class.

Hence it does not support the claim